# TuneShield: Mitigating Toxicity in Conversational AI while Fine-tuning on Untrusted Data

### v1

@@ V-A LLMs
LLaMA-2-Chat.   LLaMA-2-Chat [71] are safety-aligned LLMs from Meta, created from pre-trained LLaMA-2 models after instruction-tuning and safety alignment using RLHF. These models undergo instruction-tuning on adversarial prompts with safe responses for robustness. Safety is also integrated into the RLHF process by: (1) training a safety-specific reward model, and (2) using safety context distillation, i.e., generating safe responses with a safety pre-prompt and fine-tuning on them without it). We use LLaMA-2-Chat 7B, 13B, and 70B models for our evaluation. Note, LLaMA-2-Chat is the only safety-aligned model family in our list.
FLAN-T5.   FLAN-T5 [18] are Seq2Seq T5-based models that are instruction-tuned on 1,836 tasks across 146 categories, including chain-of-thought (CoT) reasoning tasks which claim to improve the zero-shot reasoning ability of the models. Instruction tuning also includes the toxic language detection task. We use the XXL (11B) variant.
OPT-IML.   OPT-IML [34] models are created by fine-tuning the pre-trained OPT model on a benchmark dataset containing 2K tasks. We use the 30B variants.
Vicuna.   Vicuna [17] models are instruction-tuned LLaMA-1 models [72]. Vicuna v1.3 models are created by fine-tuning 125K user conversations with ChatGPT shared on ShareGPT.com. While these models are not explicitly safety-aligned, they may have inherited some safety properties indirectly from ChatGPT. We use the 13B and 33B models.
V-B Industry API services
We use 2 popular moderation APIs from the industry to compare with our LLM-based toxicity classifiers.
@@ V-B Industry API services
Perspective API (P-API).   Perspective API [3] by Google Jigsaw is a text-based content moderation API aimed at identifying toxic content in conversations. Perspective API is built on a multilingual BERT-based model trained on data from online forums. Similar to prior work [24], we consider an input sample as toxic if the probability value for the ‘toxicity’ attribute is $\geq$ threshold of 0.5.
@@
OpenAI Moderation API (O-API).    The OpenAI moderation API [1] is built on a fine-tuned GPT model. Since the API does not provide probability values for the samples, we use on the returned ‘flagged’ attribute. An input sample is classified as toxic if the flagged attribute is returned as True. [^3] The API did not support probability outputs at the time of our study. We use the ‘flagged’ attribute as it captures several different types of unsafe content [2]. We use the ‘text-moderation-latest’ model.
V-C Toxicity classification evaluation
@@
We evaluate the performance of various LLMs in detecting toxic samples from our offensive and specialized categories and compare them with industry APIs. For offensive, we evaluate 24K samples: 12K non-toxic and 12K toxic from the non-toxic and offensive datasets, respectively. For specialized, we evaluate 4.4K samples: 2.2K non-toxic and 2.2K toxic from the non-toxic and specialized datasets, respectively.
Metrics.   We use F1-score for the toxic class to measure the performance.
Figures 5(a) and 5(b) report the F1-scores with standard deviations as error bars for the offensive and specialized categories, averaged over 10 prompt variations (Section III-C).
(a) F1-score for toxic class for the offensive category
(b) F1-score for toxic class for the specialized category
Figure 5: F1-scores for toxic class in the offensive and specialized categories
Our refusal approach, which leverage the safety properties of LLMS, significantly outperform industry API services for both offensive and specialized categories. Using the refusal approach, our best performing LLaMA-2-Chat models achieve an F1 score of 91.16% and 89.43%, compared to the best-performing industry API (O-API) achieving F1 scores of 70.97% and 78.1%, for the offensive and specialized categories, respectively. The refusal approach works best when applied to safety-aligned models. Recall that LLaMA-2-Chat is the only family in our list explicitly aligned for safety. The LLaMA-2-Chat models outperformed other models in offensive and specialized.
@@
  BB BART LLaMA-2  BB BART LLaMA-2
  RTR PPL FBD RTR PPL FBD RTR PPL FBD  RTR PPL FBD RTR PPL FBD RTR PPL FBD
Refusal  6.5 9.77 0.964 12.6 19.3 0.667 28.2 4.87 0.097    42.6 9.87 0.957 52.1 19.56 0.706 44.1 5.8 0.098
O-API  9.6 9.74 0.952 15.1 18.94 0.689 38 5.11 0.097    60.2 9.92 0.954 75.3 19.61 0.705 53 5.7 0.099
P-API  8.3 9.78 0.946 12.5 19.21 0.696 29 5.14 0.095    52.5 9.94 0.970 67.7 19.85 0.694 51.7 5.75 0.096
Table II: RTR and utility metrics for chatbots fine-tuned with TuneShield in Filter-only setting for offensive and specialized categories. "No-attack" and "Attack" rows report metrics without TuneShield applied.
@@
  BB BART LLaMA-2  BB BART LLaMA-2
@@
Table III: RTR and utility metrics for chatbots fine-tuned with TuneShield in fine-tuning with healing data (FT-Heal) setting for offensive and specialized categories. "NH" and "CH" indicate non-contextual and contextual healing, respectively.
VI TuneShield: Defense Evaluation
We conduct a rigorous evaluation of our TuneShield framework.
VI-A Evaluation setup
We incrementally evaluate different components of TuneShield, followed by an end-to-end evaluation. Recall that our defense goal is to reduce RTR to a level comparable to or lower than RTR in the no-attack setting, i.e., without any toxicity injection, with minimal degradation in model utility. We highlight the values in our tables in bold when this objective is achieved for a setting.
To implement our toxicity classifiers using the refusal approach, we select the best-performing LLaMA-2-Chat 13B (91.16% F1-score) and 7B (89.43% F1-score) models for offensive and specialized categories, respectively.
VI-B Evaluation metrics
There are two classes of metrics.
Evaluating toxicity.   The RTR metric described in Section IV is used to measure toxicity after fine-tuning.
Evaluating model utility.    We evaluate the utility of the fine-tuned chatbot after applying a defense. An effective defense should not significantly degrade utility. We aim to measure how well the model-generated responses align with the non-toxic part of the training data distribution. In addition, we also evaluate the coherence of the model-generated responses to the given input contexts. To measure utility, we generate chatbot responses for 1K benign contexts from the test set and use the following metrics:
(a) Perplexity ($\downarrow$): Perplexity (PPL) [62, 78] measures how well a chatbot predicts ground truth responses. PPL is measured as the exponent of the cross-entropy loss of a model in predicting a ground-truth response for an input. Lower PPL values indicate better model utility.
(b) Frechet Bert Distance ($\downarrow$): Frechet Bert Distance (FBD) [81] measures the distance between the generated and ground-truth response distributions and has been shown to correlated well with human judgement. The distance is computed using the mean and covariance obtained from the semantic representations extracted with a RoBERTa model (for those distributions). A lower FBD indicates that the distribution of generated responses is more similar to that of ground truth responses, and therefore means better utility.
(c) GRADE ($\uparrow$): GRADE (GRD) [33] is a reference-free metric to measure the coherence between a context and a response. GRADE uses a ConceptNet [69] model to recognize associations between words in context and response. A BERT model is used to encode both, followed by processing their concatenated outputs with an MLP layer. The values lie between 0 and 1, with a higher value indicating better coherence.
VI-C How effective is TuneShield in a filter-only setting?
We start by running TuneShield with only the filtering component. We simply filter out the toxic samples in the toxicity classification stage (Section III-B) and then use the filtered dataset to fine-tune the base model.
Table II reports the RTR and model utility metrics after applying TuneShield in the filter-only setting. Performing only filtering does not achieve our objective, i.e., after applying the defense, the RTR values do not reach the no-attack setting or below it (note the lack of entries in bold font). However, there is still a reduction in RTR for all defenses, compared to the attack setting (2nd row). In fact, our LLM classifier (refusal) outperforms (i.e., reduce RTR) both industry API services in 5 out of 6 cases. We see a minimal degradation in the model utility metrics of the chatbots with the filter-only setting. Results for the GRADE utility metric is in Table XIII (Appendix). GRADE scores also show minimal to no degradation in all cases.
Our results indicate that only using a filter does not necessarily guarantee effective toxicity mitigation. The remaining toxic samples (after filtering) can still inject toxicity into the models. This highlights the importance of integrating further steps (i.e., data healing and model alignment using DPO) beyond only using a toxicity filter.
VI-D Does fine-tuning on healing data improve toxicity mitigation?
@@
In this part, we run the TuneShield pipeline until Step 2 of the model fine-tuning stage (Section III-B), i.e., the base model is fine-tuned on the training dataset updated with the healing data. We call this the FT-Heal setting. We consider both contextual healing (CH) and non-contextual healing (NH). We do not perform model alignment using DPO in this setting. Table III shows the results after applying TuneShield in this setting. Results for GRADE utility metric for this setting are in Table XIV (Appendix). For each setting, we show results using non-contextual healing (NH) and contextual healing (CH).  [^4] To compute CH RTR, toxicity evaluation classifiers are updated on datasets reflecting the CH data distribution (Appendix -C).
@@
FT-Heal is clearly more effective than the Filter-only setting for mitigating toxicity. In Table III, for all victim models, except LLaMA-2, all defense settings successfully reduce RTR below no-attack setting. For many settings, the RTR reduces to 0%. The successful cases are highlighted in bold font. Compared to the Filter-only setting, the model utility is largely preserved, with slight degradation for PPL and FBD due to healing. Average PPL (across defense settings) for the Filter-only setting is 11.55, compared to 11.88 for the FT-Heal setting. Similarly, the average FBD for Filter-only is 0.582, compared to 0.616 for FT-Heal. We observe a similar trend for GRD as well (Table XIV in Appendix)—the average GRD score for Filter-only is 0.627, compared to 0.621 for FT-Heal.
@@
Our results show that the differences in outcomes between NH and CH are not significant. CH has a slightly higher RTR than NH. In successful cases, where the RTR of both defense settings is lower than the no-attack setting, the average RTR is 1.22% for NH vs 1.86% for CH. We observe a similar hit on PPL and FBD model utility metrics for CH vs NH. For the successful cases, we achieve an average PPL and FBD of 14.18 and 0.754 for CH, vs 12.27 and 0.690 for NH, respectively. This is a small price to pay for incorporating contextually relevant responses while mitigating toxicity.
Despite the PPL and FBD degradation, CH exhibits better dialog coherence than NH, achieving a comparable or better GRD score in 17 out of 18 cases (see Appendix Table XIV). Using CH can enable chatbots to provide more personalized and contextually relevant responses in toxic contexts, which has better usability prospects.
@@
Despite promising results for the BB and BART models in the FT-Heal setting, significant improvement is still needed in toxicity mitigation. For the LLaMA-2 victim model, multiple defense settings fail to sufficiently reduce RTR across both offensive and specialized categories. The problem is worse for the specialized category, especially with industry APIs. These results raise serious implications: (1) Toxicity mitigation failed in a widely used setting today. Compared to BB and BART, LLaMA-2 is a more widely used foundation model. Moreover, this LLaMA-2 pipeline uses the widely popular LoRA customization scheme. (2) The FT-Heal setting is insufficient when the defender uses a biased or imperfect toxicity classifier, e.g., the industry APIs in this case. We further investigate the O-API and P-API classifiers for bias issues for the specialized category and discover the following: Among the three toxic sub-categories, TA, BO and RI (see Section IV), O-API performs poorly in detecting RI samples (36.52% Recall), compared to BO (83.72% Recall) and TA (69.5% Recall). Similarly, P-API performs poorly in detecting BO and RI samples (50.51% and 19.65% Recall, respectively), compared to TA (84.0% Recall). This confirms our hypothesis that severe biases in toxicity classifiers can make toxicity mitigation more challenging, underscoring the need for our full TuneShield pipeline.
VI-E Does model alignment using DPO mitigate toxicity even when the toxicity classifier is biased?
@@
Here, we use the complete TuneShield pipeline, which includes model alignment using DPO. We focus on mitigating toxicity for the LLaMA-2 victim model, as the previous defense setting (fine-tuning with healing data) failed only for this model. The results are shown in Table IV when applying our full TuneShield pipeline. DPO-based alignment uses preference data created using the NH and CH approaches.
@@
Effectiveness of TuneShield.   Our complete TuneShield pipeline, which includes alignment using DPO as the final step, is able to effectively mitigate toxicity in almost all settings, i.e., RTR for defense is lower than the no-attack setting. The only exception is when using the O-API with CH approach. However, even for this case, the RTR is significantly lower at 15.4% compared to the RTR of 27.5% when using the FT-Heal setting. Similarly to FT-Heal setting, the CH approach exhibits better GRD scores than NH approach (5 out of 6 cases).
Computational cost of TuneShield.    Breakdown of computational costs of TuneShield are in Table XV (Appendix). After generating healing data, TuneShield requires 2.49 hours of compute time to fine-tune with model alignment, compared to 1.22 hours to fine-tune without TuneShield (when using 2x NVIDIA A100 GPUs) for the specialized category using the refusal classifier with CH approach.
DPO trade-off analysis.    We analyze trade-offs among DPO alignment parameters for LLaMA-2 chatbot for specialized category using the P-API filter and CH approach. Table XVI and XVII in Appendix show the results for DPO trade-off analysis. Recall that $\beta$ controls the divergence from the reference model; lower $\beta$ increases this divergence. Keeping LR and epochs constant, we find that decreasing $\beta$ reduces RTR and increases GRD score. However, it impacts the PPL (see Table XVI in the Appendix). Also, keeping $\beta$ and LR constant, we find that increasing the number of epochs similarly reduces RTR while increasing PPL (see Table XVII in the Appendix).
Understanding impact of biased classifiers.    We conduct a more granular analysis of the impact of biased classifiers on our different defense settings, i.e., FT-Heal vs the full TuneShield pipeline. Results are shown in Table V. Focusing on the specialized category, we present class-wise RTR for toxic classes BO, TA, and RI (Section IV). As discussed, industry APIs poorly detect RI samples, which reflects a higher RTR for RI in FT-Heal setting. However, in the full TuneShield setting, RI and other class RTRs are significantly reduced. This further validates TuneShield’s effectiveness in handling biased classifiers.
@@
Importance of using the full pipeline.   We test whether applying the model alignment step (DPO) after model fine-tuning on training dataset following the toxicity classification stage (i.e., without adding healing data as in FT-Heal) would yield effective results. We are only able to achieve an RTR of 30.87% for the specialized category (using the refusal classifier with CH data), which is still higher than RTR in the no-attack setting. In contrast, our full pipeline, which adds healing data in the model fine-tuning step, reduces RTR to 3.2%. We observe similar outcomes for the offensive category as well (results omitted for brevity).
TuneShield
@@
Table IV: RTR and utility metrics for LLaMA-2 chatbot applying TuneShield. “NH” and “CH” indicate non-contextual and contextual healing.
@@
Table V: Class-wise RTR for LLaMA-2 chatbot fine-tuned with FT-Heal and TuneShield settings for specialized, category using biased classifiers. Note that class-wise RTR values do not sum to aggregate RTR due to test set class imbalance.
VI-F What level of bias in the toxicity classifier can TuneShield tolerate?
In the previous section we saw how TuneShield can mitigate toxicity even when using biased toxicity classifiers (i.e., P-API and O-API). In this section, we further stress test TuneShield to determine how much classifier bias it can tolerate while still effectively mitigating toxicity. We artificially induce different levels of bias in the toxicity classifier by controlling the sub-category-wise Recall for the specialized category, i.e., for the TA, BO and RI toxic sub-categories. For this analysis, we evaluate the LLaMA-2 victim model.
We investigate TuneShield’s performance when classifier bias affects a single sub-category. We study a challenging setting where sub-category-wise Recall drops below 50%. We vary the Recall for a single target sub-category between 15% and 45% while keeping the default Recall (i.e., Recall for a decision threshold of 0.5) for the other two sub-categories. This approach allows us to assess TuneShield’s ability to handle filtering imperfections in individual categories of toxic language, simulating conditions often found in real-world scenarios where classifier performance might differ across various types of toxicity. Even in such challenging settings, TuneShield achieves RTR values close to or below the no-attack setting for biases that impact the BO and RI sub-categories. We note that the RTR values for the TA are still significantly lower than the setting w/o using TuneShield (i.e., RTR of ~60% as shown in Table IV). The TA category is more challenging because it is the majority sub-category in the dataset. Our results indicate that TuneShield can tolerate extreme biases in toxicity classifiers, e.g., up to 85% Recall degradation for a single toxic sub-category.
Defense setting Recall     RTR

### v2

@@ 5.1. LLM-based classifiers
LLaMA-2-Chat.   LLaMA-2-Chat (Touvron et al., 2023) are safety-aligned LLMs from Meta, obtained by instruction-tuning and RLHF on pre-trained LLaMA-2 models. They are instruction-tuned on adversarial prompts paired with safe responses for robustness. Safety in RLHF is enforced by (1) training a safety-specific reward model and (2) safety context distillation, i.e., generating safe responses with a safety pre-prompt and then fine-tuning on these responses without the pre-prompt. We evaluate LLaMA-2-Chat 7B, 13B, and 70B.
Other LLMs used for comparison.    We compare the performance of LLaMA-2-Chat with open-source instruction-tuned models from the FLAN-T5 (Chung et al., 2022), OPT-IML (Iyer et al., 2022), and Vicuna (Chiang et al., 2023) families.
5.2. Industry API services
We use 2 popular content moderation APIs from the industry.
@@ 5.2. Industry API services
Perspective API (P-API).   Perspective API (per, 2023) by Google Jigsaw is a text-based content moderation API aimed at identifying toxic content in conversations. Perspective API is built on a multilingual BERT-based model trained on data from online forums. Similar to prior work (Gehman et al., 2020), we consider an input sample as toxic if the probability value for the ‘toxicity’ attribute is $\geq$ threshold of 0.5.
@@
OpenAI Moderation API (O-API).    The OpenAI moderation API (ope, 2023) is built on a fine-tuned GPT model. Since the API does not provide probability values for the samples, we use on the returned ‘flagged’ attribute. An input sample is classified as toxic if the flagged attribute is returned as True. [^3] The API did not support probability outputs at the time of our study. We use the ‘flagged’ attribute as it captures several different types of unsafe content. We use the ‘text-moderation-latest’ model.
5.3. Toxicity classification evaluation
@@
We evaluate performance using our offensive and specialized datasets. For offensive, we evaluate 24K samples: 12K non-toxic and 12K toxic from the non-toxic and offensive datasets, respectively. For specialized, we evaluate 4.4K samples: 2.2K non-toxic and 2.2K toxic from the non-toxic and specialized datasets, respectively.
Evaluation Results.    Results are reported in Appendix Table 10. Our LLaMA-2-Chat classifiers outperform all other LLM-based classifiers and industry API services. Our best-performing LLaMA-2-Chat models achieve an average F1 score of 90.3% across the offensive and specialized categories. In contrast, industry APIs perform poorly; for example, O-API and P-API achieve only 74.5% and 70% average F1, respectively, across both toxic categories. Our results also show that LLaMA-2-Chat outperforms standard instruction-tuned models, confirming that refusal works best with safety-aligned models.
While LLM-based classifiers using our refusal approach show promising results, toxic language will keep evolving, likely degrading performance over time. For example, widely used industry APIs perform poorly on our datasets. Therefore, we design Optimus to robustly mitigate toxicity (after finetuning), even while integrating imperfect toxicity classifiers.
6. Optimus: Defense Evaluation
We conduct a rigorous evaluation of our Optimus framework.
6.1. Evaluation setup
We incrementally evaluate different components of Optimus, followed by an end-to-end evaluation. To implement our toxicity classifiers using the refusal approach, we select the best LLaMA-2-Chat 13B (91.16% F1-score) for offensive and 7B (89.43% F1-score) for specialized categories, respectively. Recall that our defense goal is to reduce RTR to a level comparable to or lower than RTR in the no-attack setting, i.e., without any toxicity injection, with minimal degradation in model utility. We highlight the values in our tables in bold when this objective is achieved for a setting.
6.2. Evaluation metrics
There are two classes of metrics.
Evaluating toxicity.   The RTR metric described in Section 4 is used to measure toxicity after fine-tuning.
Evaluating model utility.   We evaluate the utility of the fine-tuned chatbot after applying a defense. An effective defense should not significantly degrade utility. A key requirement for utility metrics is to capture how well the model-generated responses align with the non-toxic part of the training data distribution. In addition, we also evaluate the coherence of the model-generated responses with the given input contexts. To measure utility, we generate chatbot responses for 1K benign contexts from the test set and use the following metrics.
(a) Perplexity (PPL) ($\downarrow$): Perplexity  (Roller et al., 2021; Weeks et al., 2023) measures how well a chatbot predicts ground truth responses. PPL is measured as the exponent of the cross-entropy loss of a model in predicting a ground-truth response for an input. Lower PPL values indicate better model utility.
(b) Frechet Bert Distance (FBD) ($\downarrow$): Frechet Bert Distance (Xiang et al., 2021) measures the distance between the generated and ground-truth response distributions and has been shown to correlated well with human judgement. The distance is computed using the mean and covariance obtained from the semantic representations extracted with a RoBERTa model (for those distributions). A lower FBD indicates that the distribution of generated responses is more similar to that of ground truth responses, and therefore means better utility.
(c) GRADE (GRD) ($\uparrow$): GRADE (Huang et al., 2020) is a reference-free metric to measure the coherence between a context and a response. GRADE model recognizes the associations between words in context and response. A BERT model is used to encode both, followed by processing their concatenated outputs with an MLP layer. The values lie between 0 and 1, with a higher value indicating better coherence.
6.3. Can filtering alone mitigate toxicity?
We start by running Optimus with only the filtering component. We simply filter the toxic samples in the toxicity classification stage (Section 3.2) and then use the filtered dataset for fine-tuning.
@@ 6.3. Can filtering alone mitigate toxicity?
  $B$ BART LLaMA-2  $B$ BART LLaMA-2
  No-attack  4.6 2.5 8.8    13.1 10 13.1
Attack  14.2 22.8 50.8    69.9 79.8 59.6
Refusal  6.5 12.6 28.2    42.6 52.1 44.1
O-API  9.6 15.1 38    60.2 75.3 53
P-API  8.3 12.5 29    52.5 67.7 51.7
Table 1. RTR for chatbots fine-tuned with Optimus in Filter-only setting for offensive and specialized categories. ”No-attack” and ”Attack” rows report RTR without Optimus.
Table 1 reports the RTR in the filter-only setting. The RTR values do not reach the no-attack setting or below it (note the lack of bold entries) for any of the defense settings. We see minimal degradation in model utility metrics; these results are omitted for brevity.
Filtering alone does not achieve our objective. Residual toxic samples can still inject toxicity into the models. This highlights the importance of integrating further steps (i.e., data healing and model alignment using DPO) beyond a simple toxicity filter.
@@
  $B$ BART LLaMA-2  $B$ BART LLaMA-2
@@
Table 2. RTR and utility metrics for chatbots fine-tuned with Optimus in fine-tuning with healing data (FT-Heal) setting for offensive and specialized categories. ”NH” and ”CH” indicate non-contextual and contextual healing, respectively.
6.4. Can healing data improve mitigation?
@@
In this part, we run the Optimus pipeline until Step 2 of the model fine-tuning stage (Section 3.2), i.e., the base model is fine-tuned on the training dataset updated with the healing data. We call this the FT-Heal setting. We consider both contextual healing (CH) and non-contextual healing (NH). We do not perform model alignment using DPO in this setting. Table 2 shows the results for FT-Heal setting, i.e., using non-contextual healing (NH) and contextual healing (CH). Results for the GRADE utility metric for this setting are in Table 11 (Appendix).
@@
In Table 2, for all victim models, except LLaMA-2, all defense settings successfully reduce RTR below the no-attack setting (highlighted in bold font). The RTR reduces to 0% for many settings. Model utility is largely preserved, with slight degradation for PPL and FBD due to healing. Across defense settings, no-attack yields an average PPL of 10.92 and FBD of 0.564, compared to 11.88 and 0.616 for FT-Heal. We observe a similar trend for GRADE (Table 11 in Appendix), with no-attack scoring 0.628 versus 0.621 for FT-Heal.
Fine-tuning with healing data is significantly more effective than filtering alone. FT-Heal preserves model utility and reduces RTR below the no-attack baseline for all models except LLaMA-2.
@@
CH vs NH.    Our results (Table 2) show that the differences in outcomes between NH and CH are not significant. CH has a slightly higher RTR than NH. In successful cases, where the RTR of both defense settings is lower than the no-attack setting, the average RTR is 1.22% for NH vs 1.86% for CH. We observe a similar trend in utility metrics. For successful cases, CH results in an average PPL of 14.18 and FBD of 0.754, whereas NH achieves 12.27 and 0.690 respectively. However, CH offers a distinct advantage in dialog coherence, achieving a comparable or better GRADE score in 17 out of 18 cases (see Appendix Table 11). This is a small price to pay for enabling chatbots to provide personalized, contextually relevant responses while mitigating toxicity, which has better usability prospects.
@@
FT-Heal needs further improvement.    Despite promising results for the BB and BART models, multiple defense settings fail to sufficiently reduce the RTR for the LLaMA-2 model. The problem is worse for the specialized category, especially with industry APIs. These results raise serious implications: (1) Toxicity mitigation failed in a widely used setting today. Compared to BB and BART, LLaMA-2 is a more widely used foundation model. Moreover, this LLaMA-2 pipeline uses the widely popular LoRA customization scheme. (2) The FT-Heal setting is insufficient when the defender uses a biased or imperfect toxicity classifier, e.g., the industry APIs in this case.
We further investigate the O-API and P-API classifiers for bias issues for the specialized category and discover the following: Among the three toxic sub-categories, TA, BO and RI (see Section 4), O-API performs poorly in detecting RI samples (36.52% Recall), compared to BO (83.72% Recall) and TA (69.5% Recall). Similarly, P-API performs poorly in detecting BO and RI samples (50.51% and 19.65% Recall, respectively), compared to TA (84.0% Recall).
Severe biases in toxicity classifiers makes toxicity mitigation challenging, underscoring the need for our full Optimus pipeline.
6.5. Why the full pipeline matters: Mitigating toxicity despite biased classifiers
@@
We use the complete Optimus pipeline, which includes model alignment using DPO. We focus on the LLaMA-2 victim model, as the previous defense setting (FT-Heal) failed only for this model. DPO-based alignment uses preference data created using the NH and CH approaches. Table 3 shows the results.
@@
Effectiveness of Optimus.   Our complete Optimus pipeline is able to effectively mitigate toxicity in almost all settings, i.e., RTR for defense is lower than the no-attack setting. The only exception is when using the O-API with CH approach. However, even for this case, the RTR is significantly lower at 15.4% compared to the RTR of 27.5% when using the FT-Heal setting. Similarly to FT-Heal setting, the CH approach exhibits better GRD scores than NH approach (5 out of 6 cases).
@@
We also apply Optimus in the no-attack setting. We find that Optimus preserves model utility comparable to the no-attack setting and reduces RTR values lower than in the no-attack setting (6.7% and 6.4% for the offensive and specialized categories, respectively). We also evaluated applying the model alignment step (DPO) after fine-tuning following the toxicity classification stage, without adding healing data. This yielded an RTR of 30.87% for the specialized category, compared to 3.2% for our full pipeline with healing data, demonstrating the importance of using the complete pipeline (similar outcomes were observed for the offensive category, omitted for brevity).
Optimus
@@
Table 3. RTR and utility metrics for LLaMA-2 chatbot applying Optimus. “NH” and “CH” indicate non-contextual and contextual healing.
@@
Table 4. Class-wise RTR for LLaMA-2 chatbot fine-tuned with FT-Heal and Optimus settings for specialized category using biased classifiers. Note that class-wise RTR values do not sum to aggregate RTR due to test set class imbalance.
Computational cost of Optimus.    Breakdown of computational costs of Optimus are in Table 12 (Appendix). After generating healing data, Optimus requires 2.49 hours of compute time to fine-tune with model alignment, compared to 1.22 hours to fine-tune without Optimus (when using 2x NVIDIA A100 GPUs) for the specialized category using the refusal classifier with CH approach.
DPO trade-off analysis.    We analyze trade-offs among DPO alignment parameters for LLaMA-2 chatbot for the specialized category (using the P-API filter and the CH approach). Recall that $\beta$ controls the divergence from the reference model; lower $\beta$ increases this divergence. Keeping LR and epochs constant, decreasing $\beta$ reduces RTR and increases GRD score. However, it impacts the PPL (see Table 13 in the Appendix). Keeping $\beta$ and LR constant, we find that increasing the number of epochs similarly reduces RTR while increasing PPL (Table 14 in the Appendix).
Understanding impact of biased classifiers.    We conduct a more granular analysis of the impact of biased classifiers on our different defense settings, i.e., FT-Heal vs the full Optimus pipeline. Results are shown in Table 4. Focusing on the specialized category, we present class-wise RTR for toxic classes BO, TA, and RI (Section 4). As seen in the FT-Heal setting, where industry APIs poorly detect RI samples (resulting in a higher RTR for RI), the full Optimus setting significantly reduces the RTR for RI along with other classes, further validating its effectiveness against biased classifiers.
The full Optimus pipeline effectively mitigates toxicity even with biased or imperfect classifiers while preserving model utility.
6.6. What level of bias in the toxicity classifier can Optimus tolerate?
In this section, we further stress test Optimus to determine how much classifier bias it can tolerate while still effectively mitigating toxicity. We artificially induce different levels of bias in the toxicity classifier by controlling the sub-category-wise Recall for the specialized category, i.e., for the TA, BO and RI toxic sub-categories. We investigate a challenging setting where Recall drops below 50%, i.e., varying between 15%–45%) for a single target sub-category, while maintaining default Recall (i.e., Recall for a decision threshold of 0.5) for others. We evaluate the victim model, LLaMA-2. This simulates real-world variance in classifier performance.
As shown in Table 5, Optimus achieves RTR values close to or below the no-attack setting for biases impacting BO and RI sub-categories. For TA, the majority sub-category, RTR remains significantly lower than the setting w/o using Optimus, i.e., RTR of ~60% as shown in Table 3.
Optimus can tolerate extreme biases in toxicity classifiers, e.g., up to 85% Recall degradation for a single toxic sub-category.
Defense setting Recall  Overall RTR